QUESTION 4 Find the quadratic form q(X) that corresponds to the symmetric matrix

𝐴 = [ 1 2 3 2 − 1 2 3 2 − 1] ( is 3by3 matrix)

1
Expert's answer
2022-04-14T03:24:08-0400

"A=\\left[ \\begin{matrix}\t1&\t\t2&\t\t3\\\\\t2&\t\t-1&\t\t2\\\\\t3&\t\t2&\t\t-1\\\\\\end{matrix} \\right] \\\\q\\left( x \\right) ={x_1}^2-{x_2}^2-{x_3}^2+2\\cdot 2x_1x_2+2\\cdot 3x_1x_3+2\\cdot 2x_2x_3=\\\\={x_1}^2-{x_2}^2-{x_3}^2+4x_1x_2+6x_1x_3+4x_2x_3"